# AutoOPE: Automated Off-Policy Estimator Selection

### v1

@@ 3.2 Practical Implementation
However, this approach is feasible if we have access to several OPE tasks, a feature extraction function, and the ground-truth MSE. For this reason, our approach extensively rely on synthetic data. Specifically:
@@
We generate a meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ composed of several different synthetic OPE tasks. In this way, we have access to the ground-truth MSE. Also, since the objective is to obtain a model that is able to generalize in a zero-shot way, we want to train our model with a large number of OPE tasks. This would be unfeasible using real-world OPE data, which is scarce. We indeed generate 250,000 synthetic OPE tasks. Furthermore, to favor generalization of our model, we create OPE tasks with different numerical characteristics. For more details on the synthetic data generation, we refer to Appendix C.
@@
Regarding the features $g$, while theoretical results have proven a relation, for instance, with the exponentiated Rényi divergence, we want to take into account a wide selection of possible features. Hence, we designed a total of 43 features. In particular, we categorize the features into three groups: (i) Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward variance, the number of actions. (ii) Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different statistical distances and divergences. (iii) OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, or uses importance sampling. See Appendix E for their full description.
@@
The set $\mathcal{V}$ of candidate estimators include several model-free, model-based and hybrid estimators, as done in [40]. All the estimators used are described in Appendix B.
@@
As a supervised model $\hat{f}$, we use a Random Forest regressor [5] because of its high-quality predictions on tabular data and its speed, both in training and inference. For more details on training and optimization, we refer to Appendix D.
We recap the AutoOPE method in Algorithms 1 (training of the model) and 2 (zero-shot estimator selection).
@@
1:  Generate meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ (Appendix C)
@@
3:  $\hat{f}\leftarrow\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)$ (Train the random forest)
@@
The AutoOPE approach has several advantages. First of all, it is data-driven: AutoOPE tries to learn the shared patterns in several different synthetic OPE tasks with the objective to generalize to new unseen tasks. For this, we use a Random Forest, which is a powerful and fast model for tabular data. Furthermore, it is theory-inspired: it stems from theoretical findings to create a data-driven model that takes in input features such as divergences which are proven to be related to the estimator’s error, but also features that have no theoretical guarantee at the moment. Thanks to this, we can also provide insights that can shed light on future theoretical findings by analyzing the features that had a significant impact on the model’s performance. Finally, it is zero-shot. This means that a practitioner can avoid training for each new estimator selection task, leading to a much faster computational time.
A possible downside of our approach is that we trained AutoOPE on large meta-dataset $\mathcal{M}$ that is synthetic and created a priori, which may not lead to generalization to particular real-world OPE tasks. However, in Section 4, we show that AutoOPE generalizes well to unseen real-world OPE tasks with an extensive experimental evaluation on a large number of real-world OPE tasks.
@@ 4 Experiments
In this section, we describe the experimental analysis of our proposed AutoOPE and show it is competitive with the PAS-IF baseline on several real-world tasks, as well as exhibiting a lower computational cost.
@@ 4.1 Experimental Settings
Our proposed method is compared to the PAS-IF baseline, which has specific hyperparameters. One important drawback of PAS-IF is the high computational cost, see Section 4.3, which makes hyperparameter tuning impractical. [^3] We estimate that applying the same Bayesian Search used for AutoOPE would require at least a month of high-end GPU usage. For this reason we rely on the ad-hoc strategy presented in [40] which tunes the regularization parameter $\lambda$ for each new OPE task via grid search.
@@
We train the Random Forest model of AutoOPE on the synthetic meta-dataset $\mathcal{M}$ and we optimize its hyperparameters with a Bayesian search. For more details, we refer to Appendix D. In this Section, we only use the trained Random Forest in inference, in a zero-shot way (see Algorithm 2) on new unseen tasks.
@@
To evaluate AutoOPE and PAS-IF on the Estimator Selection task we use two metrics: Relative Regret and Spearman’s Rank Correlation Coefficient, as in [40], leveraging the fact that in all experiments we have access to the ground-truth MSE of each estimator. Further details are in Appendix H.
@@
Spearman’s Rank Correlation Coefficient [33] measures the similarity between two rankings, one based on the MSE predicted by the Estimator Selection method and the other based on the ground-truth MSE.
@@ 4.2 Results
Here we present the results we obtained on several public real-world datasets.
@@ 4.2.1 Open Bandit Dataset (OBD)
In this real-world experiment, we use the Open Bandit Dataset (OBD) [31], a publicly available dataset originated from a large-scale fashion e-commerce platform. The dataset encompasses three campaigns: ‘ALL’, ‘MEN’, and ‘WOMEN’. In each campaign, the deployed policy is randomly chosen at each round, selecting either the Uniform Random policy or the Bernoulli Thompson Sampling policy.
@@ 4.2.2 UCI datasets
To further strengthen our empirical evaluation of the generalization capabilities of AutoOPE, in this section we present 40 other real-world experiments conducted on 8 different datasets. In particular, we took 8 datasets from the UCI repository [1], as typically done in the OPE literature [2, 35, 6, 32]. We report an overview of these datasets in Table 2.
@@
To adapt these datasets for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features as context features and the class labels as actions, while the reward is 1 if the current action is the correct class for the given context, 0 vice versa. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using parameters $\alpha_{b}$, $\alpha_{e}$:
@@ 4.3 Computational Time Comparison
One important advantage of our AutoOPE method is that, after the pre-training phase, it is ready-to-use for inference, potentially on many different datasets and evaluation policies. On the other hand, the baseline method PAS-IF must be trained from scratch and perform hyperparameter optimization for every new evaluation policy and dataset. Due to this, the time required to perform the Estimator Selection task is very different. Figure 3 reports a comparison of the time required to perform the Estimator Selection task. As expected, AutoOPE is in the order of 10 times faster compared to PAS-IF, even though PAS-IF performs estimator selection on a GPU, while AutoOPE does it on a CPU. Further details are reported in Appendix G.
@@ 4.4 Additional Experiments
Finally, we conducted other additional experiments on more datasets, various ablation studies, as well as an analysis of the most important features of the AutoOPE model.
Feature Importance
A useful benefit of using Random Forest models is that we can compute the feature importance. For this analysis we use the Mean Decrease in Impurity (MDI), meaning that features with higher importance scores are more influential in the model decision-making process. Figure 3 shows a selection of features for the Random Forest model. The full plot is in Figure 4 of Appendix E. To assess the importance of each feature group, let us focus on the 15 most and least important features in the full plot of Figure 4 in Appendix E. The 11 policy-independent features appear to be of limited importance and correspond to 4 of the 15 least important features. Only 1 policy-dependent feature belongs to the first 15 most important features, which is the 4th moment of the logged rewards and is the second most important feature. Policy-dependent features, which are 24, are generally useful for the model, with 10 among the 15 most important features and 6 out of 24 among the 15 least important. The KL divergence is the most important feature, which suggests further theoretical study may be warranted. The 8 OPE estimator features are similarly spread, with 4 among the most important features and 5 among the least important ones. In particular, among the most important OPE estimator features there are features that characterize the type of estimator (if it uses a reward model or importance sampling or both).
The insights gained from this feature importance analysis is an important contribution. In particular, it can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
Additional Experiments and Ablation studies
We conducted additional experiments on estimator selection on synthetic datasets and on the CIFAR dataset [18]. Also, we conducted different ablation studies, investigating the impact of scaling the size of the meta-dataset, the diversity of the meta-dataset, and the features used. We describe all these additional experiments in Appendix F.
@@ 5 Conclusions
For future work, a natural direction to follow would be to try different machine learning models for AutoOPE, in particular for scenarios where one may want further guarantees on the Estimator Selection process. Furthermore, one could try to theoretically analyze the relationship between the features of the AutoOPE meta-dataset and the performance of the various estimators, which was unveiled by the feature importance of the random forest model. Finally, our approach focused on the estimator selection problem in the contextual bandit domain. Our method could be used as a stepping stone towards improving OPE estimator selection in the more general Reinforcement Learning scenario.
@@ Appendix A Related Work
In the realm of Contextual Bandit Off-Policy Evaluation, there exists a rich and diverse landscape of published research aimed at assessing the performance of policies without actually deploying them, avoiding the hazard of selecting action under not valuable or deleterious policies which, depending on the scenario, could be costly and dangerous [10, 21, 30, 17]. Indeed, this area of study has brought contributions to various application domains, including recommender systems [22], advertising [4], and precision medicine [23], reflecting its flexibility in modeling different situations. In this context, AutoOPE emerges as a novel approach, offering a unique solution based on a black-box model that leverages synthetic data and meta-learning to tackle the Estimator Selection problem.
@@
Several pioneering works have laid the foundation for evaluating counterfactual policies in the context of Contextual Bandit. Notable contributions include studies that have introduced fundamental approaches such as the Direct Method (DM) [3], Inverse Propensity Scoring (IPS) [29, 34], and Doubly Robust (DR) [38, 12, 6], which serve as the cornerstone for OPE research, since they define the three main classes of OPE estimators: model-based, model-free and hybrid.
@@
A multitude of OPE estimators, has been designed over time as extensions or combinations of DM, IPS and DR [37, 14, 35, 26, 38, 44]. These efforts collectively contribute to advance OPE methodologies.
@@
However, many estimators incorporate hyperparameters to balance the bias-variance trade-off given the Contextual Bandit environment such as [35, 26, 38, 12, 6, 44]. As reported by [32, 31] the hyperparameter tuning strongly affects the prediction quality. To address the issue of hyperparameter optimization [36] and [39] proposed SLOPE and its improved version SLOPE++, based on Lepski’s principle [20], offering improved results compared to naïve tuning procedures. Lepski’s principle allows to define an upper-bound for the Mean Squared Error without delving into the bias estimation, that is the main reason why the task of MSE estimation is challenging. SLOPE dominates all the other methods designed for hyperparameter tuning, being the state of the art technique for this task. Other existing methods like MAGIC [38], or other built-in methods of proposed estimators [35, 44] have tackled hyperparameter selection using MSE estimators, but they offer weaker guarantees compared to SLOPE.
@@
The Estimator Selection problem has emerged as a relatively unexplored but pivotal aspect of OPE research. Empirical studies by [42] and [32] have shown the fundamental role of Estimator Selection on OPE accuracy, highlighting its importance in practical applications.
@@
In this context, the recent PAS-IF (Policy-Adaptive Estimator Selection via Importance Fitting) method [40] emerges as the only baseline for Estimator Selection. PAS-IF uses a neural network to learn a subsampling rule that is used to generate two pseudo-datasets starting from the available logging data. The subsampling rule is learned by optimizing an importance fitting loss that depends on the specific logging and evaluation policies one wishes to evaluate. Once the two pseudo datasets have been generated, PAS-IF allows to estimate the candidate estimators’ Mean Squared Error based on the empirical average of rewards in these pseudo datasets.
@@
Our proposed AutoOPE method, on the other hand, competes with PAS-IF in performance by introducing a novel data-driven approach that leverages meta-learning with synthetic data and a black-box model. AutoOPE offers a versatile approach, allowing the comparison of different estimator classes and the selection of the most accurate one. This innovative approach empirically promises to significantly enhance the accuracy of OPE, reducing an important gap in the Contextual Bandit Off-Policy Evaluation field.
Appendix B Off-policy Estimators Description

### v2

@@ 3.2 Practical Implementation
In the previous section, we provided a theoretical intuition for our approach. In this section, we describe how to implement the proposed solution in practice. In particular, this approach needs several OPE tasks, a feature extraction function, and the ground-truth MSE. For this reason, we rely on synthetic data. Specifically:
@@
We generate a dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ composed of several different synthetic OPE tasks. In this way, we have access to the ground-truth MSE. Also, since the objective is to obtain a model that is able to generalize in a zero-shot way, we want to train our model with a large number of OPE tasks. This would be unfeasible using real-world OPE data, which is scarce. We indeed generate 250,000 synthetic OPE tasks. Furthermore, to favor generalization of our model, we create OPE tasks with different distributions. For more details on the synthetic data generation, we refer to Appendix D.
@@
Regarding the features $g$, while theoretical results have proven a relation, for instance, with the exponentiated Rényi divergence, we want to take into account a wide selection of possible features. Hence, we designed a total of 43 features. In particular, we categorize the features into three groups: (i) Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward variance, the number of actions. (ii) Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different statistical distances and divergences. (iii) OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, or uses importance sampling. See Appendix F for their full description.
@@
The set $\mathcal{V}$ of candidate estimators include several model-free, model-based and hybrid estimators, as done in Udagawa et al. (2023). All the estimators used are described in Appendix C.
@@
As a supervised model $\hat{f}$, we use a Random Forest regressor (Breiman, 2001) because of its high-quality predictions on tabular data and its speed, both in training and inference. For more details on training and optimization, we refer to Appendix E.
@@
1:  Generate dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ (Appendix D)
@@
3:  $\hat{f}\leftarrow\arg\min_{f}$ $\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M},\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)$ (Train the random forest)
@@
We recap the AutoOPE method in Algorithms 1 (training of the model) and 2 (zero-shot estimator selection). The AutoOPE approach has several advantages. First of all, it is data-driven: AutoOPE tries to learn the shared patterns in several different OPE tasks. Furthermore, it is theory-inspired: we create a data-driven model inspired by the fact that there are some features of the OPE tasks which are theoretically proven to be related to the estimator’s error. Finally, it is zero-shot. This means that a practitioner can avoid training for each new estimator selection task, leading to a much faster computational time.
A possible downside of our approach is that we trained AutoOPE on a synthetic data set $\mathcal{M}$ created a priori, which may not lead to generalization to particular real-world OPE tasks. In particular, Theorem 1, provides the guarantee that our method generalizes on the same distribution of tasks. This implicitly assumes that, at test time, the OPE tasks come from the same distribution of the synthetic training tasks. However, this assumption might not hold. For this reason, we relax this assumption in the following:
Theorem 2.
Let $P_{S}$ and $P_{T}$ be the distributions of synthetic and real (test) tasks respectively. Let $L_{S}(f)$ and $L_{T}(f)$ be the expected loss of $f$ with respect to $P_{S}$ and $P_{T}$ respectively. Under the same assumptions of Theorem 1, for any $\delta>0$, with probability at least $1-\delta$, we can bound $L_{T}$ of the empirical risk minimizer $\hat{f}$ as follows:
$L_{T}(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}$
where $d$ is the chi-squared divergence, defined as $d(P_{S},P_{T})=\operatorname*{\mathbb{E}}_{P_{S}}[((P_{T}(x)-P_{S}(x))/P_{S}(x))^{2}]$, and $K$ is a constant.
Also this finding is intuitive: we have that our approach learns to generalize on the test OPE tasks as long as it is trained on a large number of synthetic tasks and the divergence $d(P_{S},P_{T})$ is small. Notice that, the real OPE task distribution is unknown, and so is $d(P_{S},P_{T})$. In order to find whether $d(P_{S},P_{T})$ is small in practice or not, in Section 4, we test AutoOPE on unseen real-world OPE tasks with an extensive experimental evaluation [^3] We test on almost 50 different real-world experimental configurations, while prior work is usually tested on 2 or 3. on a large number of real-world OPE tasks, obtained from different real datasets.
@@ 4 EXPERIMENTS
In this section, we describe the experimental analysis of our proposed AutoOPE and show its strong performance compared the PAS-IF baseline on several real-world tasks, while exhibiting a lower computational cost.
@@ 4.1 Experimental Settings
Our proposed method is compared to the PAS-IF baseline, which has specific hyperparameters. One important drawback of PAS-IF is the high computational cost, see Section 4.3, which makes hyperparameter tuning impractical. [^4] We estimate that applying the same Bayesian Search used for AutoOPE would require at least a month of high-end GPU usage. For this reason we rely on the ad-hoc strategy presented in (Udagawa et al., 2023) which tunes the regularization parameter $\lambda$ for each new OPE task via grid search.
@@
We train the Random Forest model of AutoOPE on the synthetic dataset $\mathcal{M}$ and we optimize its hyperparameters with a Bayesian search. For more details, we refer to Appendix E. In this Section, we only use the trained Random Forest in inference, in a zero-shot way (see Algorithm 2) on new unseen tasks.
@@
To evaluate AutoOPE and PAS-IF on the Estimator Selection task we use two metrics: Relative Regret and Spearman’s Rank Correlation Coefficient, as in (Udagawa et al., 2023), leveraging the fact that in all experiments we have access to the ground-truth MSE of each estimator. Further details are in Appendix J.
@@
Spearman’s Rank Correlation Coefficient (Spearman, 1961) measures the similarity between two rankings, one based on the MSE predicted by the Estimator Selection method and the other based on the ground-truth MSE.
@@ 4.2.1 Open Bandit Dataset (OBD)
In this real-world experiment, we use the Open Bandit Dataset (OBD) (Saito et al., 2021a), a publicly available dataset originated from a large-scale fashion e-commerce platform. The dataset encompasses three campaigns: ‘ALL’, ‘MEN’, and ‘WOMEN’. In each campaign, the deployed policy is randomly chosen at each round, selecting either the Uniform Random policy or the Bernoulli Thompson Sampling policy.
@@ 4.2.2 UCI datasets
To further strengthen our empirical evaluation of the generalization capabilities of AutoOPE, in this section we present 40 other real-world experiments conducted on 8 different datasets. In particular, we took 8 datasets from the UCI repository (Asuncion and Newman, 2007), as typically done in the OPE literature (Beygelzimer and Langford, 2009a; Su et al., 2020a; Dudík et al., 2015; Saito et al., 2021b). We report an overview of these datasets in Table 2.
@@
To adapt these datasets for CB we perform the standard supervised to bandit conversion (Beygelzimer and Langford, 2009a; Joachims et al., 2018), reinterpreting the input features as context features and the class labels as actions, while the reward is 1 if the current action is the correct class for the given context, 0 vice versa. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using parameters $\alpha_{b}$, $\alpha_{e}$:
@@ 4.3 Computational Time Comparison
One important advantage of our AutoOPE method is that, after the pre-training phase, it is ready-to-use for inference, potentially on many different datasets and evaluation policies. On the other hand, the baseline method PAS-IF must be trained from scratch and perform hyperparameter optimization for every new evaluation policy and dataset. Due to this, the time required to perform the Estimator Selection task is very different. Figure 3 reports a comparison of the time required to perform the Estimator Selection task. As expected, AutoOPE is in the order of 10 times faster compared to PAS-IF, even though PAS-IF performs estimator selection on a GPU, while AutoOPE does it on a CPU. Further details are reported in Appendix I.
@@ 4.4 Additional Experiments
Figure 3 shows a selection of features for the Random Forest model. The full plot is in Figure 4, Appendix F, together with its discussion. The insights gained from this feature importance analysis can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance. Also, we conducted additional experiments on estimator selection on synthetic datasets and on the CIFAR dataset (Krizhevsky et al., 2009) reported in Appendix G. Finally, we conducted different ablation studies, investigating the impact of scaling the size of the synthetic dataset, the diversity of the dataset, and the features used. We describe all these ablation experiments in Appendix H.
@@ 5 CONCLUSIONS
For future work, a natural direction to follow would be to try different machine learning models for AutoOPE. Furthermore, one could try to theoretically analyze the relationship between the features of the AutoOPE dataset and the performance of the various estimators, which was unveiled by the feature importance of the random forest model. Finally, our approach focused on the estimator selection problem in the contextual bandit domain. Our method could be used as a stepping stone towards OPE estimator selection in the more general Reinforcement Learning scenario.
@@ Appendix A RELATED WORK
In the realm of Contextual Bandit Off-Policy Evaluation, there exists a rich and diverse landscape of published research aimed at assessing the performance of policies without actually deploying them, avoiding the hazard of selecting action under not valuable or deleterious policies which, depending on the scenario, could be costly and dangerous (Gilotte et al., 2018; Levine et al., 2020; Saito and Joachims, 2021; Kiyohara et al., 2021). Indeed, this area of study has brought contributions to various application domains, including recommender systems (Li et al., 2015), advertising (Bottou et al., 2013), and precision medicine (Lu et al., 2021), reflecting its flexibility in modeling different situations. In this context, AutoOPE emerges as a novel approach, offering a unique solution based on a supervised learning model that leverages synthetic data and supervised learning to tackle the Estimator Selection problem.
@@
Several pioneering works have laid the foundation for evaluating counterfactual policies in the context of Contextual Bandit. Notable contributions include studies that have introduced fundamental approaches such as the Direct Method (DM) (Beygelzimer and Langford, 2009b), Inverse Propensity Scoring (IPS) (Precup et al., 2000; Strehl et al., 2010), and Doubly Robust (DR) (Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015), which serve as the cornerstone for OPE research, since they define the three main classes of OPE estimators: model-based, model-free and hybrid.
@@
A multitude of OPE estimators, has been designed over time as extensions or combinations of DM, IPS and DR (Swaminathan and Joachims, 2015; Kallus and Uehara, 2019; Su et al., 2020a; Metelli et al., 2021; Thomas and Brunskill, 2016; Wang et al., 2017). These efforts collectively contribute to advance OPE methodologies.
@@
However, many estimators incorporate hyperparameters to balance the bias-variance trade-off given the Contextual Bandit environment such as (Su et al., 2020a; Metelli et al., 2021; Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015; Wang et al., 2017). As reported by (Saito et al., 2021b, a) the hyperparameter tuning strongly affects the prediction quality. To address the issue of hyperparameter optimization (Su et al., 2020b) and (Tucker and Lee, 2021) proposed SLOPE and its improved version SLOPE++, based on Lepski’s principle (Lepski and Spokoiny, 1997), offering improved results compared to naïve tuning procedures. Lepski’s principle allows to define an upper-bound for the Mean Squared Error without delving into the bias estimation, that is the main reason why the task of MSE estimation is challenging. SLOPE dominates all the other methods designed for hyperparameter tuning, being the state of the art technique for this task. Other existing methods like MAGIC (Thomas and Brunskill, 2016), or other built-in methods of proposed estimators (Su et al., 2020a; Wang et al., 2017) have tackled hyperparameter selection using MSE estimators, but they offer weaker guarantees compared to SLOPE.
@@
The Estimator Selection problem has emerged as a relatively unexplored but pivotal aspect of OPE research. Empirical studies by (Voloshin et al., 2021) and (Saito et al., 2021b) have shown the fundamental role of Estimator Selection on OPE accuracy, highlighting its importance in practical applications.
@@
In this context, the recent PAS-IF (Policy-Adaptive Estimator Selection via Importance Fitting) method (Udagawa et al., 2023) emerges as the only baseline for Estimator Selection. PAS-IF uses a neural network to learn a subsampling rule that is used to generate two pseudo-datasets starting from the available logging data. The subsampling rule is learned by optimizing an importance fitting loss that depends on the specific logging and evaluation policies one wishes to evaluate. Once the two pseudo datasets have been generated, PAS-IF allows to estimate the candidate estimators’ Mean Squared Error based on the empirical average of rewards in these pseudo datasets.
@@
Our proposed AutoOPE method, on the other hand, competes with PAS-IF in performance by introducing a novel data-driven approach that leverages learning from synthetic data with a supervised model. AutoOPE offers a versatile approach, allowing the comparison of different estimator classes and the selection of the most accurate one. This innovative approach empirically promises to significantly enhance the accuracy of OPE, reducing an important gap in the Contextual Bandit Off-Policy Evaluation field.
Appendix B DEFERRED PROOFS
Theorem 1.
Under Assumptions 1, 2, and 3, suppose we have $N=|\mathcal{V}||\mathcal{M}|$ independent samples $\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}=g(\hat{V}_{i},\mathcal{T}_{i})$ and $y_{i}=\text{MSE}(\hat{V}_{i}(\pi_{e});\mathcal{T}_{i})$, drawn uniformly over the estimators $\hat{V}\in\mathcal{V}$ and tasks $\mathcal{T}\in\mathcal{M}$. Let $\hat{f}$ be the empirical risk minimizer over the finite hypothesis class $\mathcal{F}$:
$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
Then, for any $\delta>0$, with probability at least $1-\delta$, the expected loss $L(\hat{f})$ satisfies:
$L(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}},$
where $\sigma^{2}$ is the subgaussian parameter from Assumption 2, and the expected loss $L(f)$ is defined as
$L(f)=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f(x)\right)\right].$
Proof.
Under Assumption 1, there exists a function $f^{*}\in\mathcal{F}$ such that for all $(\hat{V},\mathcal{T})$,
$y=\text{MSE}(\hat{V}(\pi_{e});\mathcal{T})=f^{*}(x)=f^{*}(g(\hat{V},\mathcal{T})).$
This implies that the expected loss of $f^{*}$ is zero:
$L(f^{*})=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f^{*}(x)\right)\right]=0.$
Let $\Delta(f)=L_{\mathcal{M}}(f)-L(f)$ denote the deviation between the empirical and expected losses, where
$L_{\mathcal{M}}(f)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
By Assumption 2, for any fixed $f\in\mathcal{F}$, the random variable $\mathcal{L}\left(y_{i},f(x_{i})\right)$ is subgaussian with parameter $\sigma^{2}$. Using Hoeffding’s inequality for subgaussian variables, we have for any $\epsilon>0$:
$\mathbb{P}\left(|\Delta(f)|\geq\epsilon\right)\leq 2\exp\left(-\frac{N\epsilon^{2}}{2\sigma^{2}}\right).$
Since $\mathcal{F}$ is finite, applying the union bound over all hypotheses:
$\mathbb{P}\left(\sup_{f\in\mathcal{F}}|\Delta(f)|\geq\epsilon\right)\leq 2|\mathcal{F}|\exp\left(-\frac{N\epsilon^{2}}{2\sigma^{2}}\right).$
Setting the right-hand side equal to $\delta$ and solving for $\epsilon$, we get:
$\epsilon=\sigma\sqrt{\frac{2\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}.$
On the event where $\sup_{f\in\mathcal{F}}|\Delta(f)|\leq\epsilon$, the following holds for the empirical risk minimizer $\hat{f}$:
$L(\hat{f})=L_{\mathcal{M}}(\hat{f})-\Delta(\hat{f})\leq L_{\mathcal{M}}(\hat{f})+\epsilon.$
Since $L_{\mathcal{M}}(\hat{f})\leq L_{\mathcal{M}}(f^{*})$ by definition of $\hat{f}$, and $L_{\mathcal{M}}(f^{*})=L(f^{*})+\Delta(f^{*})=0+\Delta(f^{*})\leq\epsilon$, we have:
$L(\hat{f})\leq L_{\mathcal{M}}(\hat{f})+\epsilon\leq L_{\mathcal{M}}(f^{*})+\epsilon\leq\epsilon+\epsilon=2\epsilon.$
Substituting $\epsilon$ yields the desired bound:
$L(\hat{f})\leq 2\sigma\sqrt{\frac{2\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}=\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}.$
Theorem 2.
Let $P_{S}$ and $P_{T}$ be the distribution of synthetic and real test tasks respectively. Let $L_{S}(f)$ and $L_{T}(f)$ be the expected loss of $f$ with respect to $P_{S}$ and $P_{T}$ respectively. Under the same assumptions of Theorem 1, for any $\delta>0$, with probability at least $1-\delta$, we can bound $L_{T}$ of the empirical risk minimizer $\hat{f}$ as follows:
$L_{T}(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}$
where $d$ is a divergence defined as $d(P_{S},P_{T})=\operatorname*{\mathbb{E}}_{P_{S}}[((P_{T}(x)-P_{S}(x))/P_{S}(x))^{2}]$, and $K$ is a constant.
Proof.
For any $f\in\mathcal{F}$:
$L_{T}(f)$ $=L_{S}(f)+L_{T}(f)-L_{S}(f)$ (2)
$\leq L_{S}(f)+\left\lvert L_{T}(f)-L_{S}(f)\right\rvert$ (apply absolute value)
[…]
$=L_{S}(f)+K\sqrt{d(P_{S},P_{T})}$
where $d$ is the chi-squared divergence, defined as:
[…]
If we consider the empirical risk minimizer $\hat{f}$ on $\mathcal{M}\overset{\mathrm{iid}}{\sim}P_{S}$ defined as:
$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{(x_{i},y_{i})\in\mathcal{M}}\mathcal{L}\left(y_{i},f(x_{i})\right),$
and we apply Theorem 1, we have that:
$L_{T}(\hat{f})$ $\leq L_{S}(\hat{f})+K\sqrt{d(P_{S},P_{T})}$ (applying Eq. 2)
$\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}\ ,$ (applying Theorem 1)
with probability at least $1-\delta$.
Appendix C OFF-POLICY ESTIMATORS DESCRIPTION